What would this phase look like in an actual expression? Five times a number is increased by two. This result is then decreased by three times the number. The result is three more than three times the number
Answer by ankor@dixie-net.com(22740)

What would this phase look like in an actual expression?
Five times a number is increased by two.
This result is then decreased by three times the number.
The result is three more than three times the number
:
5n + 2 - 3n = 3n + 3
combining like terms
5n - 3n - 3n = 3 - 2
5n - 6n = 1
-n = 1
mult both sides by -1
n = -1